OpenStudy (anonymous):

HELPPPPPP PLEASE ! solve each system by graphing Y=1/2x+7 Y=3/2x+3

HanAkoSolo (jamierox4ev3r):

you use the y intercept to figure out where in the y-axis the slope will intercept.

HanAkoSolo (jamierox4ev3r):

Then you use rise over run to plot the slopes

HanAkoSolo (jamierox4ev3r):

are you sure there aren't any inequalities?

OpenStudy (anonymous):

page 363

OpenStudy (anonymous):

@Jamierox4ev3r

zepdrix (zepdrix):

need help with number 11?

OpenStudy (anonymous):

do you still need help?

OpenStudy (anonymous):

yES

zepdrix (zepdrix):

\[\large y=\frac{1}{2}x+7\]\[\large y=\frac{3}{2}x+3\] Our equations are given in slope-intercept form: \[\large y=mx+b\] The \(\large b\) value is the y-intercept. It's where we cross the y-axis. So for our first equation, we see that the line crosses the y-axis at y=7.

zepdrix (zepdrix):

To draw the line, we'll need another point where it crosses. We can either find the x-intercept, or simply plug in a value for x to find another point.

zepdrix (zepdrix):

let's plug in \(\large x=2\). \(\large y=\dfrac{1}{2}(2)+7\) Which gives us, \(\large y=8\) So that gives us another cooridnate point (2,8).